# Online Variational Sequential Monte Carlo

### v1

@@ A.3. Proof of the main results
We are now ready to prove Theorem 4.6 and Corollary 4.7 of the main paper, which are restated in some more detail below. Our proofs will be based on theory developed by Karimi et al., (2019), where the minimization of a non-convex smooth objective function is considered. Thus, we assume, in Assumption A.7, that the mean field $h(\theta)$ is indeed the gradient of some smooth function of $\theta$, the so-called Lyapunov function (depending on $N$), maximized by the algorithm. However, since the ‘surrogate’ gradient considered in VSMC, whose time-normalized asymptotic limit is the target of OVSMC, is a biased approximation of the gradient of the ELBO $\mathcal{L}^{\texttt{SMC}}$ (see Section 2.2 of the main paper), the Lyapunov function does not have a straightforward interpretation in this case.
@@
(33) $\gamma_{t}\leq a\gamma_{t+1},\quad\gamma_{t}-\gamma_{t+1}\leq a^{\prime}\gamma_{t+1}^{2},\quad\gamma_{1}\leq 1/(2\tilde{\alpha}+2c_{h}),$
@@
(34) $\mathbb{E}[\lVert h(\theta_{\tau})\rVert^{2}]\leq\frac{2(d_{0,t}+c_{0,t}+(4\tilde{\alpha}^{3}+c_{\gamma})\sum_{s=0}^{t}\gamma_{s+1}^{2})}{\sum_{s=0}^{t}\gamma_{s+1}},$
@@
(35) $\mathbb{E}[\lVert h(\theta_{\tau})\rVert^{2}]=\mathcal{O}\left(\frac{\log t}{\sqrt{t}}\right),$
@@ A.4. Auxiliary results
(36) $\nabla w_{\theta}(x,f_{\theta}(x,y,v),y)=\frac{1}{r_{\theta}(x,f_{\theta}(x,y,v),y)}\Big{(}g_{\theta}(f_{\theta}(x,y,v),y)\nabla m_{\theta}(x,f_{\theta}(x,y,v))\\ +m_{\theta}(x,f_{\theta}(x,y,v))\nabla g_{\theta}(f_{\theta}(x,y,v),y)\\ -w_{\theta}(x,f_{\theta}(x,y,v),y)\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)\Big{)},$
@@
(37) $\lVert\nabla w_{\theta}(x,f_{\theta}(x,y,v),y)\rVert\leq\frac{1}{\epsilon^{2}}(\lVert\nabla m_{\theta}(x,f_{\theta}(x,y,v))\rVert+\lVert\nabla g_{\theta}(f_{\theta}(x,y,v),y)\rVert)+\frac{1}{\epsilon^{4}}\lVert\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)\rVert\\ =\frac{\kappa(2\epsilon^{2}+1)}{\epsilon^{4}}.$
@@
(38) $\left\lvert w_{\theta}(x,x^{\prime},y)-w_{\theta^{\prime}}(x,x^{\prime},y)\right\rvert\leq\frac{m_{\theta}(x,x^{\prime})\left\lvert g_{\theta}(x^{\prime},y)-g_{\theta^{\prime}}(x^{\prime},y)\right\rvert+g_{\theta^{\prime}}(x^{\prime},y)\left\lvert m_{\theta}(x,x^{\prime})-m_{\theta^{\prime}}(x,x^{\prime})\right\rvert}{r_{\theta}(x,x^{\prime},y)}\\ +m_{\theta^{\prime}}(x,x^{\prime})g_{\theta^{\prime}}(x^{\prime},y)\frac{\left\lvert r_{\theta^{\prime}}(x,x^{\prime},y)-r_{\theta}(x,x^{\prime},y)\right\rvert}{r_{\theta}(x,x^{\prime},y)r_{\theta^{\prime}}(x,x^{\prime},y)}\leq\frac{\kappa(2\epsilon^{2}+1)}{\epsilon^{4}}\lVert\theta-\theta^{\prime}\rVert.$
@@
(39) $\left\lvert w_{\theta}(x,f_{\theta}(x,y,v),y)-w_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\right\rvert\leq\frac{(2\epsilon^{2}+1)\kappa}{\epsilon^{4}}\lVert\theta-\theta^{\prime}\rVert.$
@@
(40) $\lVert\nabla w_{\theta}(x,f_{\theta}(x,y,v),y)-\nabla w_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\rVert$
(41) $\leq$ $\left\lVert\frac{g_{\theta}(f_{\theta}(x,y,v),y)\nabla m_{\theta}(x,f_{\theta}(x,y,v))}{r_{\theta}(x,f_{\theta}(x,y,v),y)}-\frac{g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))}{r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}\right\rVert$
(42) $+\left\lVert\frac{m_{\theta}(x,f_{\theta}(x,y,v))\nabla g_{\theta}(f_{\theta}(x,y,v),y)}{r_{\theta}(x,f_{\theta}(x,y,v),y)}-\frac{m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\nabla g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)}{r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}\right\rVert$
(43) $+\left\lVert\frac{w_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\nabla r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}{r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}-\frac{w_{\theta}(x,f_{\theta}(x,y,v),y)\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)}{r_{\theta}(x,f_{\theta}(x,y,v),y)}\right\rVert,$
@@
(44) $\left\lVert\frac{g_{\theta}(f_{\theta}(x,y,v),y)\nabla m_{\theta}(x,f_{\theta}(x,y,v))}{r_{\theta}(x,f_{\theta}(x,y,v),y)}-\frac{g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))}{r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}\right\rVert$
(45) $\leq$ $\frac{g_{\theta}(f_{\theta}(x,y,v),y)\lVert\nabla m_{\theta}(x,f_{\theta}(x,y,v))-\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\rVert}{r_{\theta}(x,f_{\theta}(x,y,v),y)}$
(46) $+\frac{\lVert\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\rVert\left\lvert g_{\theta}(f_{\theta}(x,y,v),y)-g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\right\rvert}{r_{\theta}(x,f_{\theta}(x,y,v),y)}$
(47) $+\lVert\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\rVert g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\frac{\left\lvert r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)-r_{\theta}(x,f_{\theta}(x,y,v),y)\right\rvert}{r_{\theta}(x,f_{\theta}(x,y,v),y)r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}$
(48) $\leq$ $\frac{\kappa}{\epsilon^{3}}(\epsilon+\epsilon^{2}\kappa+\kappa)\lVert\theta-\theta^{\prime}\rVert.$
@@
(49) $\lVert\nabla w_{\theta}(x,f_{\theta}(x,y,v),y)-\nabla w_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\rVert$
(50) $\leq\left(\frac{2\kappa}{\epsilon^{3}}(\epsilon+\epsilon^{2}\kappa+\kappa)+\frac{\kappa}{\epsilon^{5}}(\epsilon+2\epsilon^{2}\kappa+2\kappa)\right)\lVert\theta-\theta^{\prime}\rVert$
(51) $=(\epsilon+2\epsilon^{3}+2\kappa+2\epsilon^{4}\kappa+4\epsilon^{2}\kappa)\frac{\kappa}{\epsilon^{5}}\lVert\theta-\theta^{\prime}\rVert.$
@@
(52) $\tilde{\kappa}\vcentcolon=\max\left\{\frac{(2\epsilon^{2}+1)\kappa}{\epsilon^{4}},(\epsilon+2\epsilon^{3}+2\kappa+2\epsilon^{4}\kappa+4\epsilon^{2}\kappa)\frac{\kappa}{\epsilon^{5}}\right\}=(\epsilon+2\epsilon^{3}+2\kappa+2\epsilon^{4}\kappa+4\epsilon^{2}\kappa)\frac{\kappa}{\epsilon^{5}}.$
@@
(53) $|\tilde{T}_{\theta}^{t}h(z)|=|T_{\theta}^{t}h(z)-\tau_{\theta}h|\leq\varrho^{t}\|h\|_{\infty},$
where $\varrho\vcentcolon=1-\nu_{1}(\mathsf{Z})=1-\epsilon^{7N+1}$. We turn to (ii) and (iii) and introduce the short-hand notations
(54) $a_{\theta}^{k}$ :$\vcentcolon=r_{\theta}(\tilde{x}_{t-1}^{k},x_{t}^{k},y_{t}),$
(55) $\beta_{\theta}^{N}$ :$\vcentcolon=\left(\sum_{i=1}^{N}\frac{w_{\theta}(\tilde{x}_{t-1}^{i},x_{t}^{i},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\delta_{x_{t}^{i}}\right)^{\varotimes N},$
@@
(56) $\left\lvert T_{\theta}h(z_{t})-T_{\theta^{\prime}}h(z_{t})\right\rvert\leq\idotsint h(x_{t+1},y_{t+1},\tilde{x}_{t}^{1:N},v_{t+1}^{1:N})\bar{m}(x_{t},x_{t+1})\bar{g}(x_{t+1},y_{t+1})\,\mu(dx_{t+1})\,\eta(dy_{t+1})\\ \times\int_{x_{t}^{1:N}}\left\lvert\beta_{\theta}^{N}\prod_{k=1}^{N}a_{\theta}^{k}-\beta_{\theta^{\prime}}^{N}\prod_{k=1}^{N}a_{\theta^{\prime}}^{k}\right\rvert(d\tilde{x}_{t}^{1:N})\,\mu^{\varotimes N}(dx_{t}^{1:N})\,\nu^{\varotimes N}(dv_{t+1}^{1:N}).$
@@
(57) $\left\lvert\beta_{\theta}^{N}\prod_{k=1}^{N}a_{\theta}^{k}-\beta_{\theta^{\prime}}^{N}\prod_{k=1}^{N}a_{\theta^{\prime}}^{k}\right\rvert$ $\leq\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert\prod_{k=1}^{N}a_{\theta}^{k}+\left\lvert\prod_{k=1}^{N}a_{\theta}^{k}-\prod_{k=1}^{N}a_{\theta^{\prime}}^{k}\right\rvert\beta_{\theta^{\prime}}^{N}$
(58) $\leq\frac{1}{\epsilon^{N}}\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert+\left(\sum_{i^{\prime}=1}^{N}\big{|}a_{\theta}^{i^{\prime}}-a_{\theta^{\prime}}^{i^{\prime}}\big{|}\prod_{k=1}^{i^{\prime}-1}a_{\theta}^{k}\prod_{k=i^{\prime}+1}^{N}a_{\theta^{\prime}}^{k}\right)\beta_{\theta^{\prime}}^{N}$
(59) $\leq\frac{1}{\epsilon^{N}}\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert+\frac{1}{\epsilon^{N-1}}\sum_{k=1}^{N}\left\lvert a_{\theta}^{k}-a_{\theta^{\prime}}^{k}\right\rvert\beta_{\theta^{\prime}}^{N},$
@@
(60) $\sum_{k=1}^{N}\left\lvert a_{\theta}^{k}-a_{\theta^{\prime}}^{k}\right\rvert\leq N\kappa\lVert\theta-\theta^{\prime}\rVert.$
@@
(61) $\frac{1}{\epsilon^{N-1}}\sum_{k=1}^{N}\left\lvert a_{\theta}^{k}-a_{\theta^{\prime}}^{k}\right\rvert\beta_{\theta^{\prime}}^{N}\leq\frac{\kappa N}{\epsilon^{7N-1}}\lVert\theta-\theta^{\prime}\rVert\mu(\{x_{t}^{1},\ldots,x_{t}^{N}\}^{N}).$
@@
(62) $\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert=\sum_{i_{1:N}\in\{1,\ldots,N\}^{N}}\left\lvert\bar{w}_{\theta}^{i_{1:N}}-\bar{w}_{\theta^{\prime}}^{i_{1:N}}\right\rvert\delta_{x_{t}^{i_{1:N}}},$
@@
(63) $\left\lvert\bar{w}_{\theta}^{i_{1:N}}-\bar{w}_{\theta^{\prime}}^{i_{1:N}}\right\rvert$ $=\left\lvert\prod_{j=1}^{N}\frac{w_{\theta}(\tilde{x}_{t-1}^{i_{j}},x_{t}^{i_{j}},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}-\prod_{j=1}^{N}\frac{w_{\theta^{\prime}}(\tilde{x}_{t-1}^{i_{j}},x_{t}^{i_{j}},y_{t})}{\sum_{\ell=1}^{N}w_{\theta^{\prime}}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\right\rvert$
(64) $\leq\sum_{j=1}^{N}\left(\prod_{j^{\prime}=1}^{j-1}\frac{w_{\theta}(\tilde{x}_{t-1}^{i_{j^{\prime}}},x_{t}^{i_{j^{\prime}}},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\prod_{j^{\prime}=j+1}^{N}\frac{w_{\theta^{\prime}}(\tilde{x}_{t-1}^{i_{j^{\prime}}},x_{t}^{i_{j^{\prime}}},y_{t})}{\sum_{\ell=1}^{N}w_{\theta^{\prime}}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\right.$
@@
(66) $\leq\frac{1}{(\epsilon^{6}N)^{N-1}}\sum_{j=1}^{N}\left(\frac{|w_{\theta}(\tilde{x}_{t-1}^{i_{j}},x_{t}^{i_{j}},y_{t})-w_{\theta^{\prime}}(\tilde{x}_{t-1}^{i_{j}},x_{t}^{i_{j}},y_{t})|}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\right.$
@@
(68) $\leq\frac{1}{(\epsilon^{6}N)^{N-1}}\sum_{j=1}^{N}\left(\frac{\tilde{\kappa}}{N\epsilon^{3}}\lVert\theta-\theta^{\prime}\rVert+\frac{\tilde{\kappa}}{N\epsilon^{9}}\lVert\theta-\theta^{\prime}\rVert\right)$
(69) $=\frac{\tilde{\kappa}N}{N^{N}\epsilon^{6N-3}}\left(1+\frac{1}{\epsilon^{6}}\right)\lVert\theta-\theta^{\prime}\rVert,$
@@
(70) $\frac{1}{\epsilon^{N}}\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert\leq\frac{\tilde{\kappa}N}{\epsilon^{7N-3}}\left(1+\frac{1}{\epsilon^{6}}\right)\lVert\theta-\theta^{\prime}\rVert\mu(\{x_{t}^{1},\ldots,x_{t}^{N}\}^{N}).$

### v3

@@ A.3. Proof of the main results
We are now ready to prove Theorem 4.7 and Corollary 4.8, which are restated in some more detail below. Our proofs will be based on theory developed by Karimi et al., (2019), where the minimization of a non-convex smooth objective function is considered. Thus, we assume, in Assumption A.7, that the mean field $h(\theta)$ is indeed the gradient of some smooth function of $\theta$, the so-called Lyapunov function (depending on $N$), maximized by the algorithm. However, since the ‘surrogate’ gradient considered in VSMC, whose time-normalized asymptotic limit is the target of OVSMC, is a biased approximation of the gradient of the ELBO $\mathcal{L}^{\texttt{SMC}}$ (see Section 2.2), the Lyapunov function does not have a straightforward interpretation in this case.
@@
$\gamma_{t}\leq a\gamma_{t+1},\quad\gamma_{t}-\gamma_{t+1}\leq a^{\prime}\gamma_{t+1}^{2},\quad\gamma_{1}\leq 1/(2\tilde{\alpha}+2c_{h}),$ (31)
@@
$\mathbb{E}[\lVert h(\theta_{\tau})\rVert^{2}]\leq\frac{2(d_{0,t}+c_{0,t}+(4\tilde{\alpha}^{3}+c_{\gamma})\sum_{s=0}^{t}\gamma_{s+1}^{2})}{\sum_{s=0}^{t}\gamma_{s+1}},$ (32)
@@
$\mathbb{E}[\lVert h(\theta_{\tau})\rVert^{2}]=\mathcal{O}\left(\frac{\log t}{\sqrt{t}}\right),$ (33)
@@ A.4. Auxiliary results
$\nabla w_{\theta}(x,f_{\theta}(x,y,v),y)=\frac{1}{r_{\theta}(x,f_{\theta}(x,y,v),y)}\Big{(}g_{\theta}(f_{\theta}(x,y,v),y)\nabla m_{\theta}(x,f_{\theta}(x,y,v))\\ +m_{\theta}(x,f_{\theta}(x,y,v))\nabla g_{\theta}(f_{\theta}(x,y,v),y)\\ -w_{\theta}(x,f_{\theta}(x,y,v),y)\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)\Big{)},$ (34)
@@
$\lVert\nabla w_{\theta}(x,f_{\theta}(x,y,v),y)\rVert\leq\frac{1}{\epsilon^{2}}(\lVert\nabla m_{\theta}(x,f_{\theta}(x,y,v))\rVert+\lVert\nabla g_{\theta}(f_{\theta}(x,y,v),y)\rVert)+\frac{1}{\epsilon^{4}}\lVert\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)\rVert\\ =\frac{\kappa(2\epsilon^{2}+1)}{\epsilon^{4}}.$ (35)
@@
$\left\lvert w_{\theta}(x,x^{\prime},y)-w_{\theta^{\prime}}(x,x^{\prime},y)\right\rvert\leq\frac{m_{\theta}(x,x^{\prime})\left\lvert g_{\theta}(x^{\prime},y)-g_{\theta^{\prime}}(x^{\prime},y)\right\rvert+g_{\theta^{\prime}}(x^{\prime},y)\left\lvert m_{\theta}(x,x^{\prime})-m_{\theta^{\prime}}(x,x^{\prime})\right\rvert}{r_{\theta}(x,x^{\prime},y)}\\ +m_{\theta^{\prime}}(x,x^{\prime})g_{\theta^{\prime}}(x^{\prime},y)\frac{\left\lvert r_{\theta^{\prime}}(x,x^{\prime},y)-r_{\theta}(x,x^{\prime},y)\right\rvert}{r_{\theta}(x,x^{\prime},y)r_{\theta^{\prime}}(x,x^{\prime},y)}\leq\frac{\kappa(2\epsilon^{2}+1)}{\epsilon^{4}}\lVert\theta-\theta^{\prime}\rVert.$ (36)
@@
$\left\lvert w_{\theta}(x,f_{\theta}(x,y,v),y)-w_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\right\rvert\leq\frac{(2\epsilon^{2}+1)\kappa}{\epsilon^{4}}\lVert\theta-\theta^{\prime}\rVert.$ (37)
@@
$\lVert\nabla w_{\theta}(x,f_{\theta}(x,y,v),y)-\nabla w_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\rVert$ (38)
$\leq$ $\left\lVert\frac{g_{\theta}(f_{\theta}(x,y,v),y)\nabla m_{\theta}(x,f_{\theta}(x,y,v))}{r_{\theta}(x,f_{\theta}(x,y,v),y)}-\frac{g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))}{r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}\right\rVert$ (39)
$+\left\lVert\frac{m_{\theta}(x,f_{\theta}(x,y,v))\nabla g_{\theta}(f_{\theta}(x,y,v),y)}{r_{\theta}(x,f_{\theta}(x,y,v),y)}-\frac{m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\nabla g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)}{r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}\right\rVert$ (40)
$+\left\lVert\frac{w_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\nabla r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}{r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}-\frac{w_{\theta}(x,f_{\theta}(x,y,v),y)\nabla r_{\theta}(x,f_{\theta}(x,y,v),y)}{r_{\theta}(x,f_{\theta}(x,y,v),y)}\right\rVert,$ (41)
@@
$\left\lVert\frac{g_{\theta}(f_{\theta}(x,y,v),y)\nabla m_{\theta}(x,f_{\theta}(x,y,v))}{r_{\theta}(x,f_{\theta}(x,y,v),y)}-\frac{g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))}{r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}\right\rVert$ (42)
$\leq$ $\frac{g_{\theta}(f_{\theta}(x,y,v),y)\lVert\nabla m_{\theta}(x,f_{\theta}(x,y,v))-\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\rVert}{r_{\theta}(x,f_{\theta}(x,y,v),y)}$ (43)
$+\frac{\lVert\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\rVert\left\lvert g_{\theta}(f_{\theta}(x,y,v),y)-g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\right\rvert}{r_{\theta}(x,f_{\theta}(x,y,v),y)}$ (44)
$+\lVert\nabla m_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v))\rVert g_{\theta^{\prime}}(f_{\theta^{\prime}}(x,y,v),y)\frac{\left\lvert r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)-r_{\theta}(x,f_{\theta}(x,y,v),y)\right\rvert}{r_{\theta}(x,f_{\theta}(x,y,v),y)r_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)}$ (45)
$\leq$ $\frac{\kappa}{\epsilon^{3}}(\epsilon+\epsilon^{2}\kappa+\kappa)\lVert\theta-\theta^{\prime}\rVert.$ (46)
@@
$\lVert\nabla w_{\theta}(x,f_{\theta}(x,y,v),y)-\nabla w_{\theta^{\prime}}(x,f_{\theta^{\prime}}(x,y,v),y)\rVert$ (47)
$\leq\left(\frac{2\kappa}{\epsilon^{3}}(\epsilon+\epsilon^{2}\kappa+\kappa)+\frac{\kappa}{\epsilon^{5}}(\epsilon+2\epsilon^{2}\kappa+2\kappa)\right)\lVert\theta-\theta^{\prime}\rVert$ (48)
$=(\epsilon+2\epsilon^{3}+2\kappa+2\epsilon^{4}\kappa+4\epsilon^{2}\kappa)\frac{\kappa}{\epsilon^{5}}\lVert\theta-\theta^{\prime}\rVert.$ (49)
@@
$\tilde{\kappa}\vcentcolon=\max\left\{\frac{(2\epsilon^{2}+1)\kappa}{\epsilon^{4}},(\epsilon+2\epsilon^{3}+2\kappa+2\epsilon^{4}\kappa+4\epsilon^{2}\kappa)\frac{\kappa}{\epsilon^{5}}\right\}=(\epsilon+2\epsilon^{3}+2\kappa+2\epsilon^{4}\kappa+4\epsilon^{2}\kappa)\frac{\kappa}{\epsilon^{5}}.$ (50)
@@
$|\tilde{T}_{\theta}^{t}h(z)|=|T_{\theta}^{t}h(z)-\tau_{\theta}h|\leq\varrho^{t}\|h\|_{\infty},$ (51)
where $\varrho\vcentcolon=1-\nu_{1}(\mathsf{Z})=1-\epsilon^{7N+1}$.
We turn to (ii) and (iii) and introduce the short-hand notations
$a_{\theta}^{k}$ :$\vcentcolon=r_{\theta}(\tilde{x}_{t-1}^{k},x_{t}^{k},y_{t}),$ (52)
$\beta_{\theta}^{N}$ :$\vcentcolon=\left(\sum_{i=1}^{N}\frac{w_{\theta}(\tilde{x}_{t-1}^{i},x_{t}^{i},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\delta_{x_{t}^{i}}\right)^{\varotimes N},$ (53)
@@
$\left\lvert T_{\theta}h(z_{t})-T_{\theta^{\prime}}h(z_{t})\right\rvert\leq\idotsint h(x_{t+1},y_{t+1},\tilde{x}_{t}^{1:N},v_{t+1}^{1:N})\bar{m}(x_{t},x_{t+1})\bar{g}(x_{t+1},y_{t+1})\,\mu(dx_{t+1})\,\eta(dy_{t+1})\\ \times\int_{x_{t}^{1:N}}\left\lvert\beta_{\theta}^{N}\prod_{k=1}^{N}a_{\theta}^{k}-\beta_{\theta^{\prime}}^{N}\prod_{k=1}^{N}a_{\theta^{\prime}}^{k}\right\rvert(d\tilde{x}_{t}^{1:N})\,\mu^{\varotimes N}(dx_{t}^{1:N})\,\nu^{\varotimes N}(dv_{t+1}^{1:N}).$ (54)
@@
$\left\lvert\beta_{\theta}^{N}\prod_{k=1}^{N}a_{\theta}^{k}-\beta_{\theta^{\prime}}^{N}\prod_{k=1}^{N}a_{\theta^{\prime}}^{k}\right\rvert$ $\leq\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert\prod_{k=1}^{N}a_{\theta}^{k}+\left\lvert\prod_{k=1}^{N}a_{\theta}^{k}-\prod_{k=1}^{N}a_{\theta^{\prime}}^{k}\right\rvert\beta_{\theta^{\prime}}^{N}$ (55)
$\leq\frac{1}{\epsilon^{N}}\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert+\left(\sum_{i^{\prime}=1}^{N}\big{|}a_{\theta}^{i^{\prime}}-a_{\theta^{\prime}}^{i^{\prime}}\big{|}\prod_{k=1}^{i^{\prime}-1}a_{\theta}^{k}\prod_{k=i^{\prime}+1}^{N}a_{\theta^{\prime}}^{k}\right)\beta_{\theta^{\prime}}^{N}$ (56)
$\leq\frac{1}{\epsilon^{N}}\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert+\frac{1}{\epsilon^{N-1}}\sum_{k=1}^{N}\left\lvert a_{\theta}^{k}-a_{\theta^{\prime}}^{k}\right\rvert\beta_{\theta^{\prime}}^{N},$ (57)
@@
$\sum_{k=1}^{N}\left\lvert a_{\theta}^{k}-a_{\theta^{\prime}}^{k}\right\rvert\leq N\kappa\lVert\theta-\theta^{\prime}\rVert.$ (58)
@@
$\frac{1}{\epsilon^{N-1}}\sum_{k=1}^{N}\left\lvert a_{\theta}^{k}-a_{\theta^{\prime}}^{k}\right\rvert\beta_{\theta^{\prime}}^{N}\leq\frac{\kappa N}{\epsilon^{7N-1}}\lVert\theta-\theta^{\prime}\rVert\mu(\{x_{t}^{1},\ldots,x_{t}^{N}\}^{N}).$ (59)
@@
$\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert=\sum_{i_{1:N}\in\{1,\ldots,N\}^{N}}\left\lvert\bar{w}_{\theta}^{i_{1:N}}-\bar{w}_{\theta^{\prime}}^{i_{1:N}}\right\rvert\delta_{x_{t}^{i_{1:N}}},$ (60)
@@
$\left\lvert\bar{w}_{\theta}^{i_{1:N}}-\bar{w}_{\theta^{\prime}}^{i_{1:N}}\right\rvert$ $=\left\lvert\prod_{j=1}^{N}\frac{w_{\theta}(\tilde{x}_{t-1}^{i_{j}},x_{t}^{i_{j}},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}-\prod_{j=1}^{N}\frac{w_{\theta^{\prime}}(\tilde{x}_{t-1}^{i_{j}},x_{t}^{i_{j}},y_{t})}{\sum_{\ell=1}^{N}w_{\theta^{\prime}}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\right\rvert$ (61)
$\leq\sum_{j=1}^{N}\left(\prod_{j^{\prime}=1}^{j-1}\frac{w_{\theta}(\tilde{x}_{t-1}^{i_{j^{\prime}}},x_{t}^{i_{j^{\prime}}},y_{t})}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\prod_{j^{\prime}=j+1}^{N}\frac{w_{\theta^{\prime}}(\tilde{x}_{t-1}^{i_{j^{\prime}}},x_{t}^{i_{j^{\prime}}},y_{t})}{\sum_{\ell=1}^{N}w_{\theta^{\prime}}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\right.$ (62)
@@
$\leq\frac{1}{(\epsilon^{6}N)^{N-1}}\sum_{j=1}^{N}\left(\frac{|w_{\theta}(\tilde{x}_{t-1}^{i_{j}},x_{t}^{i_{j}},y_{t})-w_{\theta^{\prime}}(\tilde{x}_{t-1}^{i_{j}},x_{t}^{i_{j}},y_{t})|}{\sum_{\ell=1}^{N}w_{\theta}(\tilde{x}_{t-1}^{\ell},x_{t}^{\ell},y_{t})}\right.$ (64)
@@
$\leq\frac{1}{(\epsilon^{6}N)^{N-1}}\sum_{j=1}^{N}\left(\frac{\tilde{\kappa}}{N\epsilon^{3}}\lVert\theta-\theta^{\prime}\rVert+\frac{\tilde{\kappa}}{N\epsilon^{9}}\lVert\theta-\theta^{\prime}\rVert\right)$ (66)
$=\frac{\tilde{\kappa}N}{N^{N}\epsilon^{6N-3}}\left(1+\frac{1}{\epsilon^{6}}\right)\lVert\theta-\theta^{\prime}\rVert,$ (67)
@@
$\frac{1}{\epsilon^{N}}\left\lvert\beta_{\theta}^{N}-\beta_{\theta^{\prime}}^{N}\right\rvert\leq\frac{\tilde{\kappa}N}{\epsilon^{7N-3}}\left(1+\frac{1}{\epsilon^{6}}\right)\lVert\theta-\theta^{\prime}\rVert\mu(\{x_{t}^{1},\ldots,x_{t}^{N}\}^{N}).$ (68)